# Enhancing Data-Limited Graph Neural Networks by Actively Distilling Knowledge from Large Language Models

### v1

@@ 4.5. Graph-LLM-based Active learning
To further improve GNNs’ few-shot learning ability, we introduce a novel Graph-LLM-based active learning strategy to select valuable nodes for LLMs and add them to the training set iteratively. We seek to select $B$ nodes where GNNs exhibit low confidence in classification results, yet LLMs can offer high-quality pseudo-labels based on their inherent knowledge. Through iterative selection, we progressively enhance the GNNs’ capabilities.
@@
As indicated by the preliminary experiment results presented in Section 3, LLMs demonstrate the ability to generate high-quality pseudo-labels for nodes with higher homophily ratios and more degrees. Thus, we define an evaluation metric that combines the confidence score of GNN’s prediction, homophily ratio, and degrees to evaluate if the node in the unlabeled node set $\mathcal{V}_{u}$ is valuable for our proposed model. The evaluation metric is defined as follows:
(12) $\mathbf{S}_{GL_{i}}=\mathbf{RS}(\mathbf{p}_{i})+\mathbf{RS}(HR_{i})+\mathbf{RS}(D_{i})$
@@
where $\mathbf{S}_{GL_{i}}$ means the evaluation score for $i$-th node, $\mathbf{p}_{i}$ (i.e. $\mathbf{p}_{i}=\max(\mathbf{Z}_{i})$), $HR_{i}$, and $D_{i}$ denote the final output confidence score, the homophily ratio, and degree for $i$-th node, respectively. The $\mathbf{RS}$ represents a ranking function used to calculate scores for each evaluation metric. Specifically, we arrange the nodes in ascending order according to the evaluation metric results, excluding $\mathbf{p}_{i}$, and assign scores ranging from 0 to 1 with a step of $1/|\mathcal{V}_{u}|$. Note that $\mathbf{RS}$ assigns scores to the $\mathbf{p}_{i}$ in descending order, prioritizing nodes for which GNNs cannot generate reliable pseudo-labels.
@@
Considering the fact that some nodes can better contribute to label propagation and model improvement in the graph, we would like to add a metric to evaluate the importance of the node and to facilitate selecting the most valuable pseudo-labels. Here, we utilize neighborhood entropy reduction to assess the importance of a given node $v_{i}$ (Wang et al., 2024). Specifically, for each node $v_{i}$ in the candidate set $\mathcal{V}_{c}$, we compute the entropy reduction in the neighbors’ softmax vectors by removing $v_{i}$ from the node set $\mathcal{V}_{n}$, which contains the $v_{i}$ and its neighbors. The basic intuition is that a node is more informative when it can greatly change uncertainty (entropy) within its neighborhood. In other words, the more changes in entropy, the more important a node is. Then we rank and assign scores to these nodes based on the change of entropy. The score of entropy change is defined as follows:
(13) $\mathbf{S}_{E_{i}}=\mathbf{RS}\left(h(\hat{y}_{\mathcal{V}_{n}-v_{i}})-h(\hat{y}_{\mathcal{V}_{n}})\right)$
@@
where ${S}_{E_{i}}$ is the score of entropy change for $v_{i}$, $h(\cdot)$ denotes the entropy function, and $\hat{y}$ represents the pseudo-labels from GNNs, which is computed based on nodes’ logits vectors and the activation function. Thus, the final evaluation metric for $v_{i}$ is:
(14) $\mathbf{S}_{i}=\mathbf{S}_{GL_{i}}+\mathbf{S}_{E_{i}}$
@@
5. Experimental Results and Analysis
In this section, we present the evaluation results of our proposed few-shot node classification model on several benchmark datasets. We aim to address the following research questions:
  • RQ1: How does our proposed model perform compared with the state-of-the-art baselines under consistent settings?
@@ 5. Experimental Results and Analysis
5.1. Experimental Setup
5.1.1. Datasets
We evaluate our proposed model using three public citation datasets: Cora, Citeseer, and PubMed (Kipf and Welling, 2016). These datasets are among the most commonly utilized citation network datasets for evaluating GNN models in node classification tasks. In these datasets, nodes represent publication documents with topics serving as labels, edges depict citation links between publications, and node features are derived from the title and abstract of publications. The dataset statistics are shown in Table 2.
Table 2. Statistics of the datasets
@@ 5.1.1. Datasets
Table 3. Experimental results for Citeseer to compare with baselines, the best results are highlighted in bold
@@
      GCN       40.28±(2.66)       51.51±(3.88)       54.97±(4.04)       56.00±(5.23)
      GAT       28.70±(1.77)       34.85±(0.55)       36.42±(1.61)       37.48±(3.08)
      GraphSAGE       38.91±(2.63)       52.57±(1.57)       55.03±(2.35)       55.56±(2.9)
      MVGRL       45.74±(2.13)       53.92±(1.43)       58.47±(3.00)       58.63±(3.33)
      Meta-PN       29.67±(3.87)       42.83±(2.30)       47.49±(5.38)       35.44±(3.18)
      CGPN       31.86±(2.02)       35.89±(2.24)       39.98±(3.60)       42.15±(3.46)
      LLM-based Model       44.96±(2.19)       55.55±(2.16)       55.95±(1.81)       57.45±(2.82)
      Our Model (GAT)       33.35±(4.22)       37.37±(0.66)       38.39±(0.62)       40.20±(1.92)
      Our Model (GraphSAGE)       42.38±(3.19)       55.18±(3.99)       58.32±(4.73)       59.25±(3.94)
      Our Model (GCN)       47.35±(1.45)       57.63±(3.58)       61.50±(3.65)       63.52±(3.27)
Table 4. Experimental results for Cora to compare with baselines, the best results are highlighted in bold
@@
      GCN       44.29±(5.22)       65.90±(7.06)       70.35±(1.55)       73.60±(0.63)
      GAT       41.96±(7.18)       60.41±(7.23)       64.58±(1.76)       67.48±(0.88)
      GraphSAGE       48.12±(6.09)       61.00±(8.31)       66.29±(1.92)       71.27±(0.72)
      MVGRL       33.88±(2.14)       42.76±(6.94)       49.43±(3.05)       52.80±(3.82)
      Meta-PN       35.92±(4.63)       53.55±(2.87)       60.74±(5.49)       67.23±(2.12)
      CGPN       44.19±(5.25)       57.65±(3.25)       61.85±(2.44)       62.18±(1.72)
      LLM-based Model       51.63±(3.81)       69.90±(5.51)       72.57±(2.31)       73.61±(1.77)
      Our Model (GAT)       45.15±(6.13)       63.46±(7.97)       67.19±(2.71)       69.19±(1.71)
      Our Model (GraphSAGE)       49.69±(4.69)       67.00±(4.49)       70.83±(3.12)       74.44±(0.43)
      Our Model (GCN)       53.98±(2.15)       74.30±(2.49)       78.13±(0.33)       79.71±(1.74)
Table 5. Experimental results for PubMed to compare with baselines, the best results are highlighted in bold
@@
      GCN       56.07±(3.01)       62.74±(1.41)       64.47±(3.99)       69.20±(3.73)
      GAT       52.51±(2.66)       59.35±(0.94)       59.17±(3.42)       63.66±(3.29)
      GraphSAGE       54.03±(1.51)       59.91±(1.62)       59.40±(1.20)       65.74±(2.58)
      MVGRL       44.04±(5.58)       51.22±(1.82)       49.45±(0.28)       53.57±(1.79)
      Meta-PN       39.8±(0.92)       45.39±(4.7)       50.47±(4.65)       56.03±(4.93)
      CGPN       43.72±(4.73)       40.97±(13.13)       32.40±(13.38)       40.9±(14.22)
      LLM-based Model       57.08±(3.09)       61.54±(1.23)       68.50±(1.96)       71.31±(2.53)
      Our Model (GAT)       54.52±(3.24)       62.53±(1.56)       61.31±(2.98)       65.98±(2.40)
      Our Model (GraphSAGE)       56.13±(1.66)       63.39±(2.04)       65.21±(1.16)       70.56±(2.72)
      Our Model (GCN)       58.55±(1.03)       65.01±(0.73)       69.11±(2.61)       72.46±(3.66)
5.1.2. Implementation
@@
We randomly select n-shot (i.e. $n\times C$) samples as training data. Then, from the remaining data, we randomly choose $200$ instances as test data and $200$ as validation data, respectively. It is important to note that since we randomly select $n$-shot nodes as training nodes for the few-shot setting, the choice of seeds will influence the quality of the initial nodes, thereby impacting the classification performance. Hence, we conduct experiments with different seeds $[0,1,2]$ and use the average accuracy as our final results. For our Graph-LLM-based active learning strategies, we set the budget size $B=3$, indicating the selection of 3 samples per class in total during the graph active learning process. The balanced parameters are configured as $\alpha=0.3$ and $\beta=0.1$, and the KD temperature $\tau=3$ is used for distilling knowledge from LLMs to GNNs. Additionally, we assess the impacts of different training sizes $\mathbf{N}$, the sample size per class $B$ for active learning, the balance parameter $\alpha$ and $\beta$, and the KD temperature $\tau$ in Section 5.3.
5.1.3. Baselines
@@
We use 7 state-of-the-art GNN models as baselines, including 3 backbone GNN models: graph convolutional network (GCN) (Kipf and Welling, 2016), Graph-SAGE (Hamilton et al., 2017), and Graph Attention Network (GAT) (Veličković et al., 2018); 2 GNN based few-shot learning model: Meta-PN (Ding et al., 2022) and CGPN (Wan et al., 2021); 1 graph self-supervised model: MVGRL (Hassani and Khasahmadi, 2020); and 1 LLM-based few-shot learning method (Yu et al., 2023).
@@
  • GCN: The vanilla GCN conducts convolution operations on graph-structured data, which aggregates information from the neighboring nodes to iteratively update node representations.
@@
  • GAT: GAT incorporates an attention mechanism into GNN for feature aggregation, which allows GATs to focus on more important neighbors and get better node representations.
@@
  • Meta-PN: Meta-PN uses meta-learning and employs a bi-level optimization method to generate high-quality pseudo-labels.
@@
  • LLM-based model: This LLM-based model leverages LLMs to generate the pseudo nodes for each class, uses LM to encode these nodes, and uses an MLP to build edges.
@@
Note that the LLM-based model does not provide the original code; therefore, we independently developed the model based on the description provided in the paper. All experimental results are obtained by us under consistent settings, including the initial training nodes and initial embeddings.
5.2. Comparison with Baselines
@@
We conduct comprehensive experiments to evaluate the performance of our proposed model compared with 7 state-of-the-art baseline models under consistent settings and aim to answer RQ1. Specifically, we compare our model with seven state-of-the-art models using shots $n=[1,3,5,7]$. Additionally, we assess our model with different backbone models, including GCN, GAT, and GraphSAGE. The experiment results that are compared with baselines can be found in Table 3, Table 4, and Table 5. From the experimental results, we can observe that the LLM-based few-shot model either matches or surpasses both traditional GNN models and meta-learning-based models. Obviously, our proposed model with GCN as the backbone GNN model outperforms all state-of-the-art baselines by a large margin in different shots. For example, with the 3-shot setting, the improvement margin of accuracy is $(4,31)\%$ for Cora, $(2,23)\%$ for Citeseer, and $(2,24)\%$ for PubMed. These observations highlight that our proposed model can achieve state-of-the-art performance with fewer labeled nodes, rendering it a promising approach for few-shot node classification tasks.
From the experimental results presented in these tables, we can also observe that the backbone model influences the performance of our proposed model. Our model tends to achieve better performance when using a backbone model that already demonstrates superior performance. However, when comparing our proposed model with these backbone models, our proposed model exhibits a significant improvement in performance. This underscores the effectiveness of our approach in enhancing the learning capability of GNNs through the integration of LLMs as teacher models.
5.3. Evaluation of Hyper-parameters

### v2

@@ IV-E Graph-LLM-based Active learning
To further improve GNNs’ few-shot learning ability, we introduce a novel Graph-LLM-based active learning strategy to select valuable nodes for querying LLMs and add them to the training set iteratively. We seek to select $B$ nodes for each class where GNNs exhibit low confidence in classification results, yet LLMs can offer high-quality pseudo-labels based on their inherent knowledge. Through iterative selection, we progressively enhance the GNNs’ capabilities.
@@
As indicated by the preliminary experiment results presented in Section III, LLMs demonstrate the ability to generate high-quality pseudo-labels for nodes with higher homophily ratios and more degrees. Thus, we define an evaluation metric that combines the confidence score of GNN’s prediction, homophily ratio, and degrees to evaluate if the node in the unlabeled node set $\mathcal{V}_{u}$ is valuable for our proposed model. The evaluation metric is defined as follows:
$\mathbf{S}_{GL_{i}}=\mathbf{RS}(\mathbf{p}_{i})+\mathbf{RS}(HR_{i})+\mathbf{RS}(D_{i})$ (12)
@@
where $\mathbf{S}_{GL_{i}}$ means the evaluation score for $i$-th node, $\mathbf{p}_{i}$ (i.e. $\mathbf{p}_{i}=\max(\mathbf{Z}_{i})$), $HR_{i}$, and $D_{i}$ denote the final output confidence score, the homophily ratio, and degree for $i$-th node, respectively. The homophily ratio is calculated using labels generated from GNN. The $\mathbf{RS}$ represents a ranking function used to calculate scores for each evaluation metric. Specifically, we arrange the nodes in ascending order according to the evaluation metric results, excluding $\mathbf{p}_{i}$, and assign scores ranging from 0 to 1 with a step of $1/|\mathcal{V}_{u}|$. Note that $\mathbf{RS}$ assigns scores to the $\mathbf{p}_{i}$ in descending order, prioritizing nodes for which GNNs cannot generate reliable pseudo-labels.
@@
Considering the fact that some nodes can better contribute to label propagation and model improvement in the graph, we would like to add a metric to evaluate the importance of the node and to facilitate selecting the most valuable pseudo-labels. Here, we utilize neighborhood entropy reduction to assess the importance of a given node $v_{i}$ [44]. Specifically, for each node $v_{i}$ in the candidate set $\mathcal{V}_{c}$, we compute the entropy reduction in the neighbors’ softmax vectors by removing $v_{i}$ from the node set $\mathcal{V}_{n}$, which contains the $v_{i}$ and its neighbors. The basic intuition is that a node is more informative when it can greatly change uncertainty (entropy) within its neighborhood. In other words, the more changes in entropy, the more important a node is. Then we rank and assign scores to these nodes based on the change of entropy. The score of entropy change is defined as follows:
$\mathbf{S}_{E_{i}}=\mathbf{RS}\left(h(\hat{y}_{\mathcal{V}_{n}-v_{i}})-h(\hat{y}_{\mathcal{V}_{n}})\right)$ (13)
@@
where ${S}_{E_{i}}$ is the score of entropy change for $v_{i}$, $h(\cdot)$ is the entropy function, and $\hat{y}$ denotes the pseudo-labels from GNNs ($\hat{y}_{\mathcal{V}_{n}}$ and $\hat{y}_{\mathcal{V}_{n}-v_{i}}$ represent the pseudo labels for the nodes in the set $\mathcal{V}_{n}$ with and without $v_{i}$), which is computed based on nodes’ logits vectors and the activation function. Thus, the final evaluation metric for $v_{i}$ is:
$\mathbf{S}_{i}=\mathbf{S}_{GL_{i}}+\mathbf{S}_{E_{i}}$ (14)
@@
V Experimental Results and Analysis
In this section, we present the evaluation results of our proposed few-shot node classification model on benchmark datasets. We aim to answer the following research questions:
  • RQ1: How does our proposed model perform compared with state-of-the-art baselines under consistent settings?
@@ V Experimental Results and Analysis
V-A Experimental Setup
V-A1 Datasets
We evaluate our proposed model using three public citation datasets: Cora, Citeseer, and PubMed [6]. These datasets are among the most commonly utilized citation network datasets for evaluating GNN models in node classification tasks. In these datasets, nodes are papers with topics serving as labels. Edges depict citation links between papers, and node features are derived from the title and abstract of papers. The dataset statistics are shown in Table II.
TABLE II: Statistics of the datasets
@@ V-A1 Datasets
TABLE III: Few-shot node classification performance comparison on Citeseer. The best results are highlighted in bold
@@
      GCN       43.72±(1.22)       52.24±(1.04)       56.69±(4.39)       58.20±(3.20)
      GAT       29.02±(1.35)       33.91±(1.77)       36.33±(2.28)       38.77±(2.74)
      GraphSAGE       43.64±(2.60)       54.33±(3.20)       57.76±(3.90)       58.11±(3.36)
      MVGRL       46.13±(3.05)       55.61±(2.14)       59.56±(2.56)       60.52±(3.06)
      Meta-PN       31.00±(4.89)       42.84±(4.70)       47.83±(4.30)       52.33±(3.33)
      CGPN       34.75±(2.79)       41.69±(1.81)       45.88±(3.24)       46.80±(3.32)
      LLM-based Model       44.63±(1.72)       55.75±(0.63)       58.65±(1.22)       59.53±(2.39)
      Our Model (GAT, GPT)       32.86±(2.29)       39.51±(1.37)       41.55±(2.05)       42.87±(1.84)
      Our Model (GraphSAGE, GPT)       45.41±(3.70)       56.63±(3.43)       59.87±(3.73)       60.87±(3.49)
      Our Model (GCN, LLama)       45.11±(1.82)       56.22±(2.59)       60.23±(3.89)       62.05±(2.64)
      Our Model (GCN, Gemini)       47.04±(2.75)       57.01±(3.07)       60.02±(3.19)       62.55±(3.48)
      Our Model (GCN, GPT)       47.74±(1.69)       57.43±(3.33)       62.19±(2.80)       63.32±(2.64)
TABLE IV: Few-shot node classification performance comparison on Core. The best results are highlighted in bold
@@
      GCN       48.63±(5.49)       66.27±(5.04)       72.15±(2.24)       74.98±(0.97)
      GAT       41.14±(7.24)       62.07±(6.96)       65.27±(3.40)       68.26±(1.29)
      GraphSAGE       49.81±(4.66)       64.94±(6.89)       69.80±(2.41)       73.37±(1.19)
      MVGRL       29.02±(2.49)       39.60±(3.17)       41.46±(4.11)       47.34±(6.01)
      Meta-PN       36.59±(4.66)       54.55±(3.37)       59.74±(4.58)       66.89±(2.66)
      CGPN       47.69±(6.97)       55.39±(5.01)       60.57±(2.05)       62.62±(2.86)
      LLM-based Model       48.96±(3.57)       68.45±(4.80)       72.65±(1.92)       74.36±(1.35)
      Our Model (GAT, GPT)       47.40±(2.49)       66.66±(3.02)       68.47±(3.26)       70.97±(1.77)
      Our Model (GraphSAGE, GPT)       51.55±(2.15)       68.23±(3.76)       72.54±(2.40)       75.39±(0.44)
      Our Model (GCN, LLama)       52.57±(1.72)       73.07±(2.23)       73.12±(2.84)       77.26±(2.04)
      Our Model (GCN, Gemini)       53.35±(3.68)       73.42±(1.83)       75.44±(1.03)       79.61±(1.02)
      Our Model (GCN, GPT)       53.82±(2.26)       74.32±(1.79)       78.13±(1.63)       79.73±(1.08)
TABLE V: Few-shot node classification performance comparison on PubMed. The best results are highlighted in bold
@@
      GCN       54.55±(3.06)       62.64±(0.85)       66.53±(3.25)       69.99±(1.31)
      GAT       49.53±(2.87)       59.35±(0.80)       59.95±(2.22)       64.58±(1.74)
      GraphSAGE       54.62±(0.02)       60.58±(2.82)       63.79±(1.53)       68.15±(1.67)
      MVGRL       41.54±(7.91)       51.96±(4.74)       52.14±(4.35)       54.56±(2.19)
      Meta-PN       40.13±(0.50)       45.72±(4.38)       51.47±(3.84)       56.70±(5.47)
      CGPN       42.73±(3.48)       39.79±(5.86)       41.67±(5.74)       50.02±(3.33)
      LLM-based Model       56.45±(0.74)       61.89±(2.47)       67.47±(4.30)       71.25±(2.01)
      Our Model (GAT, GPT)       52.77±(3.21)       62.06±(1.47)       63.41±(0.88)       67.78±(2.95)
      Our Model (GraphSAGE, GPT)       57.37±(1.66)       63.40±(2.71)       65.98±(3.36)       71.00±(1.55)
      Our Model (GCN, LLama)       57.11±(1.16)       63.16±(1.25)       68.00±(3.00)       73.00±(3.74)
      Our Model (GCN, Gemini)       57.29±(2.22)       64.01±(0.85)       69.24±(2.93)       73.05±(2.56)
      Our Model (GCN, GPT)       58.17±(1.13)       65.40±(0.84)       69.37±(2.47)       73.43±(3.36)
V-A2 Implementation
@@
Following the traditional dataset split setting, we divide the dataset into three parts: $60\%$ for training, $20\%$ for validation, and $20\%$ for testing. From the training set, we then randomly select $n$-shot samples (i.e. $n\times C$) to be used as training data. It is important to note that since we randomly select $n$-shot nodes as training nodes for the few-shot setting, the choice of seeds will influence the quality of the initial nodes, thereby impacting the classification performance. Hence, we conduct experiments with different seeds $[0,1,2]$ and use the average accuracy as our final results. For our Graph-LLM-based active learning strategies, we set the budget size $B=3$, indicating the selection of 3 samples per class in total during the graph active learning process. The balanced parameters are configured as $\alpha=0.3$ and $\beta=0.1$, and the KD temperature $\tau=3$ is used for distilling knowledge from LLMs to GNNs. We use GPT3.5-Turbo as our base LLM model. Additionally, we assess the impacts of different backbone models (GCN, GAT, and GraphSAGE), different LLM base models (GPT3.5-Turbo, Gemini-1.5-flash, and LLama-3.1-8b), different training sizes $\mathbf{N}$, the sample size per class $B$ for active learning, the balance parameter $\alpha$ and $\beta$, and the KD temperature $\tau$ in V-B and V-C. Note that for the LLM base models, we use the API for GPT and Gemini, while LLama is run locally.
V-A3 Baselines
@@
We use 7 state-of-the-art models as baselines: 3 backbone GNN models (GCN [6], GAT [26], and GraphSAGE [8]), 2 GNN based few-shot learning models (Meta-PN [14], CGPN [15]), 1 graph self-supervised model (MVGRL [45]), and 1 LLM-based few-shot learning model [25].
@@
  • GCN: GCN conducts convolution operations on graph-structured data, which aggregates information from the neighbors to iteratively update node representations.
@@
  • GAT: GAT incorporates an attention mechanism into GNN for feature aggregation, which allows GAT to focus on more important neighbors and get better node representations.
@@
  • Meta-PN: Meta-PN uses meta-learning and employs a bi-level optimization to generate high-quality pseudo-labels.
@@
  • LLM-based model: This LLM-based model leverages LLMs to generate the pseudo nodes for each class, uses LM to encode these nodes and uses an MLP to build edges.
@@
Note that the LLM-based model does not provide the original code; therefore, we independently developed the model based on the paper. For fair comparison, the hyperparameters of all the models are tuned on the validation set. All experimental results are conducted under consistent settings.
V-B Comparison with Baselines
@@
We conduct comprehensive experiments to evaluate the performance of our proposed model compared with 7 state-of-the-art baseline models under consistent settings and aim to answer RQ1. Specifically, we compare our model with seven state-of-the-art models using shots $n=[1,3,5,7]$. Additionally, we assess our model with different backbone models (GCN, GAT, GraphSAGE) and LLMs (GPT, Gemini, LLama). The experiment results can be found in Table IV, Table III, and Table V. From the experimental results, we can observe that the LLM-based few-shot model either matches or surpasses both traditional GNN models and meta-learning-based models. Obviously, our proposed model with GCN outperforms all state-of-the-art baselines by a large margin in different shots. For example, with the 3-shot setting, the improvement margin of accuracy is $(6\sim 35)\%$ for Cora, $(2\sim 23)\%$ for Citeseer, and $(5\sim 20)\%$ for PubMed. Compared to the LLM-based few-shot model, our model performs better while requiring fewer detailed outputs, thus reducing costs and resource usage. These observations highlight that our proposed model can achieve state-of-the-art performance with fewer labeled nodes, rendering it a promising approach for few-shot node classification tasks.
From the tables, we can also observe that both the backbone models and LLMs influence the performance of our proposed model. Our model tends to perform better when paired with a high-performing backbone model. Similarly, if the LLM excels in text inference tasks, our model’s performance improves accordingly. Despite this dependency, our model significantly outperforms the backbone models alone, highlighting its ability to effectively distill knowledge from LLMs by leveraging their enhanced rationales and soft labels with few labeled data.
V-C Hyper-parameter Sensitivity Analysis